# Competition of $\chi_{c}(2P)$ quarkonia and continuum in $e^{+}e^{-}\to e^{+}e^{-}D\bar{D}$

### v1

@@ I Introduction
The ground state properties of $P$-wave charmonia $\chi_{cJ}$ ($J=0,1,2$) are relatively well known. Their decay widths and branching fractions were measured [1]. In contrast, the excited $P$-wave charmonia (their masses, widths, and branching fractions) are poorly known. Recently the production of axial $\chi_{c1}(2P)$ state in $pp$ [2] and $e^{+}e^{-}$ [3] collisions was discussed. These studies indicate that also for the famous $X(3872)$, although potentially an exotic state, the $c\bar{c}$ component can play an important role. This naturally raises the question of whether similar situation can occur for the $J^{P}=0^{+}$ and $2^{+}$ states. A number of observed states are considered as candidates for the first excited $P$-wave charmonia $\chi_{c0,2}(2P)$. A candidate for the $\chi_{c2}(2P)$ state ($\chi_{c2}(3930)$ in PDG [1]) was observed by the Belle Collaboration [4] and confirmed by the BaBar Collaboration [5] in the two-photon fusion process $e^{+}e^{-}\to e^{+}e^{-}(\gamma\gamma\to D\bar{D})$ as a resonance in the $D^{+}D^{-}$ and $D^{0}\bar{D}^{0}$ invariant mass distributions. In the $D^{0}{\bar{D}}^{0}$ channel, one can also observe a broad enhancement at $M\sim 3.8$ GeV. The intriguing problem of searching for $\chi_{c0}(2P)$ was a topic of several works, see [6, 7]. The candidate for the $\chi_{c0}(2P)$ state is either the $X(3860)$ ($\chi_{c0}(3860)$ in PDG) or the $X(3915)$ ($\chi_{c0}(3915)$ in PDG). The $X(3915)$ state is relatively narrow ($\Gamma\simeq 20$ MeV) and was observed in two-photon fusion in the $\omega J/\psi$ decay mode [8] as well as in the $B$ meson decays, $B\to KJ\psi\omega$ (Belle [9], BaBar [10]) and $B^{+}\to K^{+}D^{+}D^{-}$ (LHCb [11]). The observation of a broad resonance identified as a new charmonium-like state $X(3860)$ with preference for $J^{PC}=0^{++}$ over $2^{++}$ was reported by the Belle Collaboration [12] in the process $e^{+}e^{-}\to J/\psi D\bar{D}$. In line with the arguments presented there, the $X(3860)$ state seems to be a better candidate for the $\chi_{c0}(2P)$ charmonium state than the $X(3915)$. There is still much uncertainty about the above assignments for the $\chi_{c0,2}(2P)$ states due to limited information on their production and decay properties, and therefore further confirmation is required.
@@
The nature of the $\chi_{c0}(3915)$ and $\chi_{c2}(3930)$ charmonium-like states is also not well understood. In Ref. [13] the authors suggest that the two resonances $X(3915)$ and $X(3930)$ can be regarded as the same $J^{PC}=2^{++}$ state. It was shown in [14], using the framework of a constituent quark model, that this may be related to the molecular structure for these two states. Further considerations on the $X(3915)/X(3930)$ puzzle can be found in [15] and in the references therein.
@@ II Production of $D\bar{D}$ pairs in $e^{+}e^{-}$ collisions via $\gamma\gamma$ fusion
In the present studies, we perform calculations for the considered exclusive $2\to 4$ reaction, see the diagrams shown in Fig. 1. The fully differential cross section reads
@@ II.1 Resonant production: $\gamma^{*}\gamma^{*}\to\chi_{c2}\to D\bar{D}$
The amplitude for the $\gamma^{*}(q_{1})\gamma^{*}(q_{2})\to\chi_{c2}\to D(p_{3})\bar{D}(p_{4}))$ subprocess is given by
@@ II.1.1 Two photons vertex: $\gamma^{*}\gamma^{*}\to\chi_{c2}$
where $\hat{s}=(q_{1}+q_{2})^{2}=2(q_{1}\cdot q_{2})+q_{1}^{2}+q_{2}^{2}$. We have $q_{1}\cdot q_{2}=(M_{\chi_{c2}}^{2}+Q_{1}^{2}+Q_{2}^{2})/2$. Introducing an extra factor $(q_{1}\cdot q_{2})/M_{\chi_{c0}}$ the transition form factor $\tilde{F}_{TT,2}(Q_{1}^{2},Q_{2}^{2})$ remains dimensionless. For the vertex function (2.17) we find the relations
@@
In order to get compatibility with the formula (2.7), we write
@@ II.1.2 $\chi_{c2}\to D\bar{D}$ vertex
We noted that the most significant partial widths for $\chi_{c2}(2P)$ are $\Gamma(\chi_{c2}\to D\bar{D})$ and $\Gamma(\chi_{c2}\to D\bar{D}^{*})$. Different calculations in the literature support the fact that the total width of $\chi_{c2}(2P)$ is largely saturated by the $D\bar{D}$ mode. For instance, in Eq. (24) and Fig. 24 of [32] one can find $\Gamma(\chi_{c2}\to D\bar{D})\approx 26(12)$ MeV and $\Gamma(\chi_{c2}\to D\bar{D}^{*})\approx 22(14)$ MeV. In Table I of [33] one can find $\Gamma(\chi_{c2}\to D\bar{D})=24$ MeV and $\Gamma(\chi_{c2}\to D\bar{D}^{*})=13.6-16.6$ MeV (depending on model). In Table IV of [34], the following values corresponding to $M_{\chi_{c2}(2P)}=3930$ MeV are stated $\Gamma(\chi_{c2}\to D\bar{D})=20.4$ MeV and $\Gamma(\chi_{c2}\to D\bar{D}^{*})=6.9$ MeV. On the other hand, the authors of [14] found a different decomposition of partial widths.
@@ II.2 Resonant production: $\gamma^{*}\gamma^{*}\to\chi_{c0}\to D\bar{D}$
The narrower total width of $\chi_{c0}(3860)$ seems to be consistent with BaBar data for $D^{+}D^{-}$ channel; see Fig. 3(b) below. Similar values are given in Table VI of [35] (see $2^{3}P_{0}$ state), in Table III of [33], and in Sec. III A and Table II of [34].
The amplitude for the $\gamma^{*}(q_{1})\gamma^{*}(q_{2})\to\chi_{c0}\to D(p_{3})\bar{D}(p_{4}))$ subprocess is given by
@@
In Table II of [34] one can find $\Gamma(\chi_{c0}\to D\bar{D})=21.0(16.4)$ MeV for $M_{\chi_{c0}(2P)}=3836.8(3862.0)$ MeV. Taking $M_{\chi_{c0}}=3862$ MeV, $\Gamma(\chi_{c0}\to D\bar{D})=16.4$ MeV, and using Eqs. (2.33), (2.34) we obtain $|g_{\chi_{c0}D\bar{D}}|=2.52$.
@@ II.3 Continuum production of neutral $D$ mesons: $\gamma^{*}\gamma^{*}\to D^{0}\bar{D}^{0}$
$+i\Gamma^{(D^{*0}D^{0}\gamma)}_{\kappa_{1}\mu}(-\hat{p}_{u},q_{1})\,i\tilde{\Delta}^{(D^{*0})\,\kappa_{1}\kappa_{2}}(s_{34},\hat{p}_{u}^{2})\,i\Gamma^{(D^{*0}D^{0}\gamma)}_{\kappa_{2}\nu}(\hat{p}_{u},q_{2}).$
@@
$F^{(D^{*0})}(\hat{p}^{2})$ $=$ $\exp\left(\frac{\hat{p}^{2}-m_{D^{*0}}^{2}}{\Lambda_{D^{*0}}^{2}}\right)\,,\qquad\Lambda_{D^{*0}}=3.5\;{\rm GeV}\,$, (2.38)
@@
Here $\hat{p}^{2}<0$ and $q^{2}<0$. We take $\Lambda_{D^{*0}}=3.5\;{\rm GeV}$ estimated from a comparison of the model results to the experimental data (see Fig. 3).
The coupling constant $g_{D^{*0}D^{0}\gamma}$ is related to the decay width of $D^{*0}(2007)^{0}\to D^{0}\gamma$ as follows
@@
We use in (2.35) $\tilde{\Delta}^{(D^{*0})\,\kappa_{1}\kappa_{2}}(s_{34},\hat{p}^{2})$ the propagators for the reggeized vector meson $D^{*0}(2007)^{0}$. Two scenarios for the vector-meson reggeization effect were considered in [41, 42]. In the present work, we proceed analogously to (3.20)–(3.22) of [41]. The $D^{*0}$-meson propagator and the transverse function reads
@@ III Results and discussion
The experimental cross section for $\chi_{c2}(3930)$ production determined by the BaBar Collaboration (Eq. (17) of [5]) is
@@ III.1 $e^{+}e^{-}\to e^{+}e^{-}\chi_{c2}(3930)$
A comparison of our $2\to 3$ cross sections to the BaBar result (3.2), $\Gamma_{\gamma\gamma}\times B(\chi_{c2}\to D\bar{D})=0.24\pm 0.06~{\rm keV}$, favours rather smaller values for $B(\chi_{c2}\to D\bar{D})$.
Taking into account the condition $p_{t,\chi_{c2}}<0.05$ GeV we obtain $\sigma(e^{+}e^{-}\to e^{+}e^{-}\chi_{c2})=0.625$ pb (Cornell) and $0.405$ pb (BT). These results with comparison to (3.2) leads to $B(\chi_{c2}\to D\bar{D})=1.19$ and 1.83 for Cornell and BT potential, respectively. This suggests that the value given in (3.1) does not include cuts.
@@ III.2 $e^{+}e^{-}\to e^{+}e^{-}D\bar{D}$
Figure 2: Distributions in $\log_{10}(p_{t1}/1\,\rm{GeV})$, $|t_{1}|$, and $p_{t,D\bar{D}}$ for the $e^{+}e^{-}\to e^{+}e^{-}(\chi_{c2}(3930)\to D^{0}{\bar{D}}^{0})$ reaction for $\sqrt{s}=10.54$ GeV. We show the results without and with the cut on $p_{t,{\rm sum}}<0.05$ GeV. The calculations were done for the BT potential and with $g_{\chi_{c2}D\bar{D}}=10.1$.
We compare our theoretical results for the $e^{+}e^{-}\to e^{+}e^{-}D\bar{D}$ reaction calculated for $\sqrt{s}=10.54$ GeV with the Belle data from Fig. 2 of [4] and BaBar data from Fig. 5 of [5]. The $\chi_{c0}(3860)$ and $\chi_{c2}(3930)$ contributions are calculated for the BT potential with $R^{\prime}_{2P}(0)=0.326$ GeV5/2 (see Tables 1 and 2) and with $g_{\chi_{c2}D\bar{D}}=10.1$ (see Table 3) and $g_{\chi_{c0}D\bar{D}}=2.52$. In Fig. 3(a)-(c) we show the invariant mass distributions for the combined $D\bar{D}$ final state and for the $D^{0}\bar{D}^{0}$ and $D^{+}D^{-}$ channels separately. We can see that the complete model results, containing the continuum and resonant contributions, can reasonably well describe the experimental data. There is a significant interference effect between the $\chi_{c0}(3860)$ resonance and continuum terms. The $D^{0}\bar{D}^{0}$-continuum contribution increases from the reaction threshold, reaches a maximum at $M_{D^{0}\bar{D}^{0}}\sim 3.8$ GeV, and has a large share in the $\chi_{c2}(3930)$ resonance region. From the results, it is clearly visible that the $D^{+}D^{-}$-continuum contribution is small. Since the form factor $F(\hat{p}_{t}^{2},\hat{p}_{u}^{2},p_{34}^{2})$ is set to 1 in the calculation, our estimate should be viewed as an upper limit, especially at large $M_{D^{+}D^{-}}$. In Fig. 3(d) we show the distribution in $p_{t,D\bar{D}}$ in the region $M_{D\bar{D}}\in(3.91,3.95)$ GeV.
@@
Finally, in Fig. 4 we present our results obtained for (2.24) for the $\chi_{c0}(3860)$ resonance. In this case, we have that the increase in $\Gamma_{\chi_{c0}}$ implies a worse description of low-$M_{D\bar{D}}$ data. The contributions from the $\chi_{c2}(3930)$ resonance and $D^{0}\bar{D}^{0}$-continuum are dominant. The term from $\chi_{c0}(3860)$ is found to be small but the interference effect between the $\chi_{c0}(3860)$ and the $D^{0}\bar{D}^{0}$ continuum is significant.

### v2

@@ I Introduction
The ground state properties of $P$-wave charmonia $\chi_{cJ}$ ($J=0,1,2$) are relatively well known. Their decay widths and branching fractions were measured [1]. In contrast, the excited $P$-wave charmonia (their masses, widths, and branching fractions) are poorly known. Recently, the production of axial $\chi_{c1}(2P)$ state in $pp$ [2] and $e^{+}e^{-}$ [3] collisions was discussed. These studies indicate that also for the famous $X(3872)$, although potentially an exotic state, the $c\bar{c}$ component can play an important role. This naturally raises the question of whether similar situation can occur for the $J^{P}=0^{+}$ and $2^{+}$ states. A number of observed states are considered as candidates for the first excited $P$-wave charmonia $\chi_{c0,2}(2P)$. A candidate for the $\chi_{c2}(2P)$ state ($\chi_{c2}(3930)$ in PDG [1]) was observed by the Belle Collaboration [4] and later confirmed by the BaBar Collaboration [5] in the two-photon fusion process $e^{+}e^{-}\to e^{+}e^{-}(\gamma\gamma\to D\bar{D})$ as a resonance in the $D^{+}D^{-}$ and $D^{0}\bar{D}^{0}$ invariant mass distributions. In the $D^{0}{\bar{D}}^{0}$ channel, one can also observe a broad enhancement at $M\sim 3.8$ GeV. The intriguing problem of searching for $\chi_{c0}(2P)$ was a topic of several works; see [6, 7]. The candidate for the $\chi_{c0}(2P)$ state is either the $X(3860)$ ($\chi_{c0}(3860)$ in PDG) or the $X(3915)$ ($\chi_{c0}(3915)$ in PDG). The $X(3915)$ state is relatively narrow ($\Gamma\simeq 20$ MeV) and was observed in two-photon fusion in the $\omega J/\psi$ decay mode [8] as well as in the $B$ meson decays, $B\to KJ/\psi\omega$ (Belle [9], BaBar [10]) and $B^{+}\to D^{+}D^{-}K^{+}$ (LHCb [11]). The observation of a broad resonance identified as a new charmonium-like state $X(3860)$ with a preference for $J^{PC}=0^{++}$ over $2^{++}$ was reported by the Belle Collaboration [12] in the process $e^{+}e^{-}\to J/\psi D\bar{D}$. In line with the arguments presented there, the $X(3860)$ state seems to be a better candidate for the $\chi_{c0}(2P)$ charmonium state than the $X(3915)$. There is still much uncertainty about the above assignments for the $\chi_{c0,2}(2P)$ states due to limited information on their production and decay properties, and therefore further confirmation is required.
@@
The nature of the $\chi_{c0}(3915)$ and $\chi_{c2}(3930)$ charmonium-like states is also not well understood. In Ref. [13] the authors suggest that the two resonances $X(3915)$ and $X(3930)$ can be regarded as the same $J^{PC}=2^{++}$ state. It was shown in [14], using the framework of a constituent quark model, that this may be related to the molecular structure for these two states. Further considerations on the $X(3915)/X(3930)$ puzzle can be found in [15] and in the references therein. Prompt hadroproduction of the $\chi_{c2}(3930)$ resonance was observed by the LHCb Collaboration in the decay modes $\chi_{c2}(3930)\to D^{+}D^{-}$ and $D^{0}\bar{D}^{0}$ in proton-proton collisions [16]. A more recent LHCb amplitude analysis of the process $B^{+}\to D^{+}D^{-}K^{+}$ finds distinct $J^{PC}=0^{++}$ and $2^{++}$ states decaying to $D^{+}D^{-}$ [11], which can be identified as the $\chi_{c0}(3915)$ and $\chi_{c2}(3930)$ charmonium states, respectively.
@@ II Production of $D\bar{D}$ pairs in $e^{+}e^{-}$ collisions via $\gamma\gamma$ fusion
In the present studies, we perform calculations for the considered exclusive $2\to 4$ reaction; see the diagrams shown in Fig. 1. The fully differential cross section reads
@@ II.1 Resonant production: $\gamma^{*}\gamma^{*}\to\chi_{c2}\to D\bar{D}$
The amplitude for the $\gamma^{*}(q_{1})\gamma^{*}(q_{2})\to\chi_{c2}\to D(p_{3})\bar{D}(p_{4})$ subprocess is given by
@@ II.1.1 Two photons vertex: $\gamma^{*}\gamma^{*}\to\chi_{c2}$
where $\hat{s}=(q_{1}+q_{2})^{2}=2(q_{1}\cdot q_{2})+q_{1}^{2}+q_{2}^{2}$. We have $q_{1}\cdot q_{2}=(M_{\chi_{c2}}^{2}+Q_{1}^{2}+Q_{2}^{2})/2$. Introducing an extra factor $(q_{1}\cdot q_{2})/M_{\chi_{c2}}$ the transition form factor $\tilde{F}_{TT,2}(Q_{1}^{2},Q_{2}^{2})$ remains dimensionless. For the vertex function (2.17) we find the relations
@@
In order to get compatibility with the formula given by Eq. (2.7), we write
@@ II.1.2 $\chi_{c2}\to D\bar{D}$ vertex
We noted that the most significant partial widths for $\chi_{c2}(2P)$ are $\Gamma(\chi_{c2}\to D\bar{D})$ and $\Gamma(\chi_{c2}\to D\bar{D}^{*})$. Different calculations in the literature support the fact that the total width of $\chi_{c2}(2P)$ is largely saturated by the $D\bar{D}$ mode. For instance, in Eq. (24) and Fig. 24 of [33] one can find $\Gamma(\chi_{c2}\to D\bar{D})\approx 26(12)$ MeV and $\Gamma(\chi_{c2}\to D\bar{D}^{*})\approx 22(14)$ MeV. In Table I of [34] one can find $\Gamma(\chi_{c2}\to D\bar{D})=24$ MeV and $\Gamma(\chi_{c2}\to D\bar{D}^{*})=13.6-16.6$ MeV (depending on model). In Table IV of [35], the following values corresponding to $M_{\chi_{c2}(2P)}=3930$ MeV are stated $\Gamma(\chi_{c2}\to D\bar{D})=20.4$ MeV and $\Gamma(\chi_{c2}\to D\bar{D}^{*})=6.9$ MeV. On the other hand, the authors of [14, 37] found a different decomposition of partial widths.
@@ II.2 Resonant production: $\gamma^{*}\gamma^{*}\to\chi_{c0}\to D\bar{D}$
The narrower total width of $\chi_{c0}(3860)$ seems to be consistent with BaBar data for $D^{+}D^{-}$ channel; see Fig. 3(b) below. Similar value is given in Table VI of [36] (see $2^{3}P_{0}$ state).
The amplitude for the $\gamma^{*}(q_{1})\gamma^{*}(q_{2})\to\chi_{c0}\to D(p_{3})\bar{D}(p_{4})$ subprocess is given by
@@
There is considerable uncertainty in the literature regarding the estimation of open-charm strong decay width. For instance, in Sec. III A and Table II of [35] one can find $\Gamma(\chi_{c0}\to D\bar{D})=21.0(16.4)$ MeV for $M_{\chi_{c0}}=3836.8(3862.0)$ MeV, which is rather narrow. Similar values for the open-charm strong decay width are predicted in [34] within the linear (screened) potential model, and the results are: $\Gamma(\chi_{c0}\to D\bar{D})=22(28)$ MeV for $M_{\chi_{c0}}=3869(3848)$ MeV. In the analysis of [37] the width is predicted to be 16.6 MeV, which is in agreement with the results of [34, 35] where the adopted decay model is also the $^{3}P_{0}$ pair creation model. In [38], a smaller width about 11 MeV was obtained, while the analysis of [39] in a $^{3}P_{0}$ model obtained a large width of 110–180 MeV [which is consistent with (2.24)]. The recent analysis of [40] predicts the strong decay width of 73 MeV for $\chi_{c0}(2P)$ state. This suggest that further precise measurement of the $\chi_{c0}(3860)$ state parameters and more theoretical studies of the $\chi_{c0}(2P)$ state are still needed.
Taking $M_{\chi_{c0}}=3862$ MeV, $\Gamma(\chi_{c0}\to D\bar{D})=16.4$ MeV, and using Eqs. (2.33), (2.34) we obtain $|g_{\chi_{c0}D\bar{D}}|=2.52$. This value for $g_{\chi_{c0}D\bar{D}}$ is used in our calculations presented in Sec. III. Considering the Buchmüller-Tye potential from Table 2 and assuming the branching fraction $B(\chi_{c0}\to D\bar{D})=\Gamma(\chi_{c0}\to D\bar{D})/\Gamma_{\chi_{c0}}=16.4~{\rm MeV}/50.0~{\rm MeV}=0.328$ we get
$\Gamma_{\gamma\gamma}\times B(\chi_{c0}\to D\bar{D})=2.171~{\rm keV}\times 0.328=0.71~{\rm keV}\,.$ (2.35)
It should be emphasized that the estimate of $\Gamma_{\gamma\gamma}\times B(\chi_{c0}\to D\bar{D})$ is highly uncertain in view of the uncertainty of $\Gamma_{\gamma\gamma}$ (dependent on the choice of potential model) and $B(\chi_{c0}\to D\bar{D})$ regarding the choice of total width $\Gamma_{\chi_{c0}}$, (2.24) or (2.25), and of $\Gamma(\chi_{c0}\to D\bar{D})$ (see the discussion above). Note that the present experimental widths are not well determined. Clearly, more investigations are necessary to confirm the property of the $\chi_{c0}(2P)$.
@@ II.3 Continuum production of neutral $D$ mesons: $\gamma^{*}\gamma^{*}\to D^{0}\bar{D}^{0}$
$+\,i\Gamma^{(D^{*0}D^{0}\gamma)}_{\kappa_{1}\mu}(-\hat{p}_{u},q_{1})\,i\tilde{\Delta}^{(D^{*0})\,\kappa_{1}\kappa_{2}}(s_{34},\hat{p}_{u}^{2})\,i\Gamma^{(D^{*0}D^{0}\gamma)}_{\kappa_{2}\nu}(\hat{p}_{u},q_{2})\,$,
@@
$F^{(D^{*0})}(\hat{p}^{2})$ $=$ $\exp\left(\frac{\hat{p}^{2}-m_{D^{*0}}^{2}}{\Lambda_{D^{*0}}^{2}}\right)\,$, (2.39)
@@
Here $\hat{p}^{2}<0$ and $q^{2}<0$. We take $\Lambda_{D^{*0}}=3.3-3.5\;{\rm GeV}$ estimated from a comparison of the model results to the experimental data (see Figs. 35).
The coupling constant $g_{D^{*0}D^{0}\gamma}$ is related to the decay width of $D^{*}(2007)^{0}\to D^{0}\gamma$ as follows
@@
We use in (2.36) $\tilde{\Delta}^{(D^{*0})\,\kappa_{1}\kappa_{2}}(s_{34},\hat{p}^{2})$ the propagators for the reggeized vector meson $D^{*}(2007)^{0}$. Two scenarios for the vector-meson reggeization effect were considered in [46, 47]. In the present work, we proceed analogously to (3.20)–(3.22) of [46]. The $D^{*0}$-meson propagator and the transverse function reads
@@ III Results and discussion
The experimental cross section (efficiency corrected) for $\chi_{c2}(3930)$ production, determined by the BaBar Collaboration (Eq. (17) of [5]), is
@@ III.1 $e^{+}e^{-}\to e^{+}e^{-}\chi_{c2}(3930)$
Comparing the results (3.3) to the BaBar result (3.2) ($\Gamma_{\gamma\gamma}\times B(\chi_{c2}(2P)\to D\bar{D})=0.24\pm 0.06~{\rm keV}$) one can see that they are roughly consistent within the uncertainty.
Taking into account the condition $p_{t,\chi_{c2}}<0.05$ GeV we obtain $\sigma(e^{+}e^{-}\to e^{+}e^{-}\chi_{c2}(2P))=0.625$ pb (Cornell) and $0.405$ pb (BT). These results with comparison to (3.2) leads to $B(\chi_{c2}\to D\bar{D})=1.19$ and 1.83 for Cornell and BT potential, respectively. This suggests that the value given in (3.1) does not include cuts.
@@ III.2 $e^{+}e^{-}\to e^{+}e^{-}D\bar{D}$
From Table 3 one can see that the obtained value of the branching fraction $B(\chi_{c2}(2P)\to D\bar{D})$ depends on the potential model (the two-photon width of the $\chi_{c2}(3930)$ state) assumed in the calculations. For the results calculated without kinematical cuts, the obtained value for the product $\Gamma_{\gamma\gamma}\times B(\chi_{c2}(2P)\to D\bar{D})$ is consistent with existing Belle and BaBar measurements within uncertainties. From our analysis, the corresponding branching fraction is $B(\chi_{c2}(3930)\to D\bar{D})=0.38\pm 0.08$ for the Cornell potential and $0.58\pm 0.13$ for the Buchmüller-Tye potential. The latter value is compatible with the one reported in [51].
@@
Figure 2: Distributions in $p_{t,D\bar{D}}$, $\log_{10}(p_{t1}/1\,\rm{GeV})$ and $|t_{1}|$ for the $e^{+}e^{-}\to e^{+}e^{-}(\chi_{c2}(3930)\to D^{0}{\bar{D}}^{0})$ reaction for $\sqrt{s}=10.54$ GeV. We show the results without and with the cut on $p_{t,{\rm sum}}<0.05$ GeV. The calculations were done for the BT potential and with $g_{\chi_{c2}D\bar{D}}=10.1$.
We compare our theoretical results for the $e^{+}e^{-}\to e^{+}e^{-}D\bar{D}$ reaction calculated for $\sqrt{s}=10.54$ GeV with the Belle data from Fig. 2 of [4] and BaBar data from Fig. 5 of [5]. The $\chi_{c0}(3860)$ and $\chi_{c2}(3930)$ contributions are calculated for the BT potential with $R^{\prime}_{2P}(0)=0.326$ GeV5/2 (see Tables 1 and 2) and with $g_{\chi_{c2}D\bar{D}}=10.1$ (see Table 3), and $g_{\chi_{c0}D\bar{D}}=2.52$. In Fig. 3(a)-(c) we show the invariant mass distributions for the combined $D\bar{D}$ final state and for the $D^{0}\bar{D}^{0}$ and $D^{+}D^{-}$ channels separately. We can see that the complete model results, containing the continuum and resonant contributions, can reasonably well describe the experimental data. There is a significant interference effect between the $\chi_{c0}(3860)$ resonance and continuum terms. The $D^{0}\bar{D}^{0}$-continuum contribution increases from the reaction threshold, reaches a maximum at $M_{D^{0}\bar{D}^{0}}\sim 3.8$ GeV, and has a large share in the $\chi_{c2}(3930)$ resonance region. In these calculations, for the $D^{0}\bar{D}^{0}$-continuum contribution, we use $\Lambda_{D^{*0}}=3.5~{\rm GeV}$ in (2.39). From the presented results, it is clearly visible that the $D^{+}D^{-}$-continuum contribution is small. Since the form factor $F(\hat{p}_{t}^{2},\hat{p}_{u}^{2},p_{34}^{2})$ is set to 1 in the calculation, our estimate should be viewed as an upper limit, especially at large $M_{D^{+}D^{-}}$. In Fig. 3(d) we show the distribution in $p_{t,D\bar{D}}$ in the region $M_{D\bar{D}}\in(3.91,3.95)$ GeV.
@@
For illustrative purposes, in Fig. 4 we present our results obtained for $\Gamma_{\chi_{c0}}=200$ MeV (2.24) for the $\chi_{c0}(3860)$ resonance. In this case, we have that the increase in $\Gamma_{\chi_{c0}}$ implies a worse description of low-$M_{D\bar{D}}$ data (especially for the $D^{+}D^{-}$ BaBar data). Here, we keep the coupling constants the same as in the previous scenario. In general, using the larger total width, the coupling constant of $\chi_{c0}\to D\bar{D}$ could be consistently re-estimated. If a larger value for $B(\chi_{c0}\to D\bar{D})$ is used, this can be compensated for by a smaller two-photon width (see Table 2). From the right panel of Fig. 4 we see that the contributions from the $\chi_{c2}(3930)$ resonance and $D^{0}\bar{D}^{0}$-continuum are dominant. The term from $\chi_{c0}(3860)$ is found to be small but the interference effect between the $\chi_{c0}(3860)$ and the $D^{0}\bar{D}^{0}$ continuum is significant.
@@
An additional comment is in order here. We are not sure whether the Belle data shown in Fig. 3 was corrected for experimental efficiencies. If not, this could partly explain why our results are higher than the Belle data. Therefore, we have also compared our model results to the efficiency-corrected BaBar data, which was provided only for the combined $D\bar{D}$ final state.
Finally, in Fig. 5 we show our complete model results including $D^{0}\bar{D}^{0}$ continuum, $D^{+}D^{-}$ continuum, $\chi_{c0}(3860)$, and $\chi_{c2}(3930)$ contributions together with the efficiency-corrected (combined) BaBar data taken from Fig. 10 of [5]. In the calculation we used the Buchmüller-Tye potential for the excited $2P$ states, $\chi_{c0,2}(2P)$. The $D^{0}\bar{D}^{0}$-continuum contribution from $t/u$-channel vector-meson $D^{*}(2007)^{0}$ exchanges depends on the parameter of the off-shell form-factor (2.39). In the calculations we take two values of the form-factor parameter $\Lambda_{D^{*0}}=3.3~{\rm GeV}$ and $3.5~{\rm GeV}$ corresponding to the lower and upper line in the bands shown in Fig. 5, respectively. The presented calculations correspond to two scenarios of the $\chi_{c0}(3860)$ total width: (1) $\Gamma_{\chi_{c0}}=50$ MeV with $\chi^{2}/{\rm dof}=2.44$ (solid upper line) and $\chi^{2}/{\rm dof}=1.61$ (solid lower line); (2) $\Gamma_{\chi_{c0}}=200$ MeV with $\chi^{2}/{\rm dof}=2.61$ (dashed upper line) and $\chi^{2}/{\rm dof}=2.06$ (dashed lower line). We also checked that limiting to the range of $2m_{D}\leq M_{D\bar{D}}<3.9~{\rm GeV}$ the resulting values of $\chi^{2}/{\rm dof}$ are 1.82–2.04 and 1.39–2.85 for scenarios (1) and (2), respectively.
Figure 5: Comparison of our complete model results to the efficiency-corrected (combined) BaBar data from Fig. 10 of [5]. The black solid lines represent results obtained for $\Gamma_{\chi_{c0}}=50$ MeV (2.25), while the blue dashed lines correspond to $\Gamma_{\chi_{c0}}=200$ MeV (2.24). Here, the calculations for the $D^{0}\bar{D}^{0}$-continuum contribution were made for $\Lambda_{D^{*0}}=3.3$ and $3.5~{\rm GeV}$ in (2.39), which correspond to the lower and upper lines in the bands, respectively. We used the Buchmüller-Tye potential for the excited $2P$ states, $\chi_{c0,2}(2P)$.